### v4

# Shortcuts and Identifiability in Concept-based Models from a Neuro-Symbolic Lens

Abstract
Concept-based Models are neural networks that learn a concept extractor to map inputs to high-level concepts and an inference layer to translate these into predictions. Ensuring these modules produce interpretable concepts and behave reliably in out-of-distribution is crucial, yet the conditions for achieving this remain unclear. We study this problem by establishing a novel connection between Concept-based Models and reasoning shortcuts (RSs), a common issue where models achieve high accuracy by learning low-quality concepts, even when the inference layer is fixed and provided upfront. Specifically, we extend RSs to the more complex setting of Concept-based Models and derive theoretical conditions for identifying both the concepts and the inference layer. Our empirical results highlight the impact of RSs and show that existing methods, even combined with multiple natural mitigation strategies, often fail to meet these conditions in practice.
1 Introduction
Concept-based Models (CBMs) are a broad class of self-explainable classifiers [alvarez2018towards, 10, 39, 21, 54, 78, 63] designed for high performance and ante-hoc interpretability. Learning a CBM involves solving two conjoint problems: acquiring high-level concepts describing the input (e.g., an image) and an inference layer that predicts a label from them. In many applications, it is essential that these two elements are “high quality”, in the sense that: i) the concepts should be interpretable, as failure in doing so compromises understanding [71, 62] and steerability [teso2023leveraging, 28], both key selling points of CBMs; and ii) the concepts and inference layer should behave well also out of distribution (OOD), e.g., they should not pick up spurious correlations between the input, the concepts and the output [geirhos2020shortcut, bahadori2021debiasing, stammer2021right].
This raises the question of when CBMs can acquire “high-quality” concepts and inference layers. While existing studies focus on concept quality [49, 99, 55], they neglect the role of the inference layer altogether. In contrast, we cast the question in terms of whether it is possible to identify from data concepts and inference layers with the intended semantics, defined formally in Section˜3.1. We proceed to answer this question by building a novel connection with reasoning shortcuts (RSs), a well-known issue in Neuro-Symbolic (NeSy) AI whereby models achieve high accuracy by learning low-quality concepts even if the inference layer is fixed [9, 56, 84, 91]. For instance, a NeSy model for autonomous driving whose inference layer encodes the traffic laws might confuse pedestrians with red lights, as both entail the same prediction (the car has to stop) [4]. We generalize RSs to CBMs in which the inference layer is learned, and concept supervision may be absent. Our analysis shows that shortcuts can be exponentially many, even more than RSs, (we count them explicitly in Section˜9.5) and that maximum likelihood training is insufficient for attaining intended semantics. This hinders both interpretability and OOD behavior. On the positive side, we also specify conditions under which (under suitable assumptions) CBMs cannot be fooled by RSs, proving that the ground-truth concepts and inference layer can be identified (see Theorem˜3.9).
Our evaluation on several learning tasks suggest that CBMs can be severely impacted by reasoning shortcuts in practice, as expected, and also that the benefits of popular mitigation strategies do not carry over to this more challenging problem. These results cast doubts on the ability of these models to identify concepts and inference layers with the intended semantics unless appropriately nudged.
Contributions: In summary, we: (i) generalize reasoning shortcuts to the challenging case of CBMs whose inference layer is learned (Section˜3.1), (ii) study conditions under which maximum likelihood training can identify good concepts and inference layers (Section˜3.2), and (iii) present empirical evidence that well-tested mitigations fail in this challenging setting (Section˜5).
Figure 1: Joint reasoning shortcuts. The goal is to predict whether the sum of two MNIST digits is odd (as in Example˜2.1) from a training set of all possible unique (even, even), (odd, odd), and (odd, even) pairs of MNIST digits. Green elements are fixed, purple ones are learned. Left: ground-truth concepts and inference layer. Middle: NeSy-CBMs with given knowledge can learn reasoning shortcuts, i.e., concepts with unintended semantics. Right: CBMs can learn joint reasoning shortcuts, i.e., both concepts and inference layer have unintended semantics.
2 Preliminaries
Concept-based Models (CBMs) first map the input $\bm{\mathrm{x}}\in\mathbb{R}^{n}$ into $k$ discrete categorical concepts $\bm{\mathrm{c}}=(c_{1},\ldots,c_{k})\in\mathcal{C}$ via a neural backbone, and then infer labels $\bm{\mathrm{y}}\in\mathcal{Y}$ from this using a white-box layer, e.g., a linear layer. This setup makes it easy to figure out what concepts are most responsible for any prediction, yielding a form of ante-hoc concept-based interpretability. Several architectures follow this recipe, including approaches for converting black-box neural networks into CBMs [96, 86, 19, 52].
A key issue is how to ensure the concepts are interpretable. Some CBMs rely on concept annotations [39, 69, 21, 54, 36, 16]. These are however expensive to obtain, prompting researchers to replace them with (potentially unreliable [31, 77]) annotations obtained from foundation models [61, 93, 74, 15] or unsupervised concept discovery [alvarez2018towards, 10, 78, 70].
Neuro-Symbolic CBMs (NeSy-CBMs) specialize CBMs to tasks in which the prediction $\bm{\mathrm{y}}\in\mathcal{Y}$ ought to comply with known safety or structural constraints. These are supplied as a formal specification – a logic formula $\mathsf{K}$, aka knowledge – tying together the prediction $\bm{\mathrm{y}}$ and the concepts $\bm{\mathrm{c}}$. In NeSy-CBMs, the inference step is a symbolic reasoning layer that steers [18, 20, 89] or guarantees [47, 50, 24, 30, 1] the labels and concepts to be logically consistent according to $\mathsf{K}$. Throughout, we will consider this example task:
Example 2.1 (MNIST-SumParity).
Given two MNIST digits [43], we wish to predict whether their sum is even or odd. The numerical values of the two digits can be modelled as concepts $\bm{\mathrm{C}}\in\{0,\ldots,9\}^{2}$, and the inference layer is entirely determined by the prior knowledge: $\mathsf{K}=((y=1)\Leftrightarrow\text{$(C_{1}+C_{2})$ is odd})$. This specifies that the label $y\in\{0,1\}$ ought to be consistent with the predicted concepts $\bm{\mathrm{C}}$. See Fig.˜1 for an illustration.
Like CBMs, NeSy-CBMs are usually trained via maximum likelihood and gradient descent, but without concept supervision. The reasoning layer is typically imbued with fuzzy [97] or probabilistic [14] logic semantics to ensure differentiability. Many NeSy-CBMs require $\mathsf{K}$ to be provided upfront, as in Example˜2.1, hence their inference layer has no learnable parameters. Starting with Section˜3, we will instead consider NeSy-CBMs that – just like regular CBMs – learn the inference layer [85, 48, 13, 79, 87].
2.1 Reasoning Shortcuts
Before discussing reasoning shortcuts, we need to establish a clear relationship between concepts, inputs, and labels. The RS literature does so by assuming the following data generation process [56, 91, 80]: each input $\bm{\mathrm{x}}\in\mathbb{R}^{n}$ is the result of sampling $k$ ground-truth concepts $\bm{\mathrm{g}}=(g_{1},\ldots,g_{k})\in\mathcal{G}$ (e.g., in MNIST-SumParity two numerical digits) from an unobserved distribution $p^{*}(\bm{\mathrm{G}})$ and then $\bm{\mathrm{x}}$ itself (e.g., two corresponding MNIST images) from the conditional distribution $p^{*}(\bm{\mathrm{X}}\mid\bm{\mathrm{g}})$. [^1] This distribution subsumes stylistic factors, e.g., calligraphy. Labels $\bm{\mathrm{y}}\in\mathcal{Y}$ are sampled from the conditional distribution of $\bm{\mathrm{g}}$ given by $p^{*}(\bm{\mathrm{Y}}\mid\bm{\mathrm{g}};\mathsf{K})$ consistently with $\mathsf{K}$ (e.g., $y=1$ if and only if $g_{1}+g_{2}$ is odd). The ground-truth distribution is thus:
$p^{*}(\bm{\mathrm{X}},\bm{\mathrm{Y}})=\mathbb{E}_{\bm{\mathrm{g}}\sim p^{*}(\bm{\mathrm{G}})}[p^{*}(\bm{\mathrm{X}}\mid\bm{\mathrm{g}})p^{*}(\bm{\mathrm{Y}}\mid\bm{\mathrm{g}};\mathsf{K})]$ (1)
Intuitively, a reasoning shortcut (RS) occurs when a NeSy-CBM with fixed knowledge $\mathsf{K}$ attains high or even perfect label accuracy by learning concepts $\bm{\mathrm{C}}$ that differ from the ground-truth ones $\bm{\mathrm{G}}$.
Example 2.2.
In MNIST-SumParity, a NeSy model can achieve perfect accuracy by mapping each pair of MNIST images $\bm{\mathrm{x}}=(\bm{\mathrm{x}}_{1},\bm{\mathrm{x}}_{2})$ to the corresponding ground-truth digits, that is, $\bm{\mathrm{c}}=(g_{1},g_{2})$. However, it would achieve the same accuracy if it were to map it to $\bm{\mathrm{c}}=(g_{1}\ \mathrm{mod}\ 2,g_{2}\ \mathrm{mod}\ 2)$, as doing so leaves the parity of the sum unchanged, see Fig.˜1. Hence, a NeSy model cannot distinguish between the two based on label likelihood alone during training.
RSs by definition yield good labels in-distribution, yet they compromise out-of-distribution (OOD) performance. For instance, in autonomous driving tasks, NeSy-CBMs can confuse the concepts of “pedestrian” and “red light”, leading to poor decisions for OOD decisions where the distinction matters [4]. The meaning of concepts affected by RSs is unclear, affecting understanding [71], intervenability [72, 22, 76], debugging [lertvittayakumjorn2020find, stammer2021right] and down-stream applications that hinge on concepts being high-quality, like NeSy formal verification [88, 98, 59]. Unfortunately, existing works on RSs do not apply to CBMs and NeSy-CBMs where the inference layer is learned.
As commonly done, we work in the setting with equal discrete predicted $\mathcal{C}$ and ground-truth $\mathcal{G}$ concept sets, i.e., $\mathcal{G}=\mathcal{C}$ [56, 91]. Notice that, we make no assumption on how the set $\mathcal{G}$ is made; multiple concept vocabularies at different levels of abstraction may be valid for a given task. At this stage, different choices of $\mathcal{G}$ are allowed but may lead to distinct results for RSs, depending on the number of ground-truth concepts and how they are related to the labels. Both these two aspects will be made clear in light of the data generation process as per Assumptions˜3.1 and 3.2.
Figure 2: Examples of semantics in MNIST-SumParity restricted to $\bm{\mathrm{g}},\bm{\mathrm{c}}\in\{0,1,2\}^{2}$ for readability. Left: ideally, $\bm{\mathrm{\alpha}}$ should be the identity (i.e., $\bm{\mathrm{C}}$ recovers the ground-truth concepts $\bm{\mathrm{G}}$) and the inference layer should learn $\bm{\mathrm{\beta}}^{*}$. Middle: $(\bm{\mathrm{\alpha}}_{\mathrm{IS}},\bm{\mathrm{\beta}}_{\mathrm{IS}})\neq(\mathrm{id},\bm{\mathrm{\beta}}^{*})$ has intended semantics (Definition˜3.3), i.e., the ground-truth concepts and inference layer can be recovered and generalize OOD. Right: $(\bm{\mathrm{\alpha}}_{\mathrm{JRS}},\bm{\mathrm{\beta}}_{\mathrm{JRS}})$ affected by the Joint Reasoning Shortcuts in Fig.˜1. Elements (predicted concepts $\bm{\mathrm{C}}$ and entries in $\bm{\mathrm{\beta}}$) in red are never predicted nor used, highlighting simplicity bias. Maps are visualized as matrices.
3 Reasoning Shortcuts in CBMs
Given a finite set $\mathcal{S}$, we indicate with $\Delta_{\mathcal{S}}\subset[0,1]^{|\mathcal{S}|}$ the simplex of probability distributions $P(Q)$ over items in $\mathcal{S}$. Any random variable $Q\in\mathcal{S}$ defines a point in the simplex via its distribution $P(Q)$. Notice that the set of the simplex vertices $\mathsf{Vert}(\Delta_{\mathcal{S}})$ contains all point mass distributions $\mathbbm{1}\!\{Q=q\}$ for all $q\in\mathcal{S}$. All relevant notation we will use is reported in Table˜6.
CBMs as pairs of functions. CBMs and NeSy-CBMs differ in how they implement the inference layer, hence to bridge them we employ the following unified formalism. Any CBM can be viewed as a pair of learnable functions implementing the concept extractor and the inference layer, respectively, cf. Fig.˜2. Formally, the former is a function $\bm{\mathrm{f}}:\mathbb{R}^{n}\to\Delta_{\mathcal{C}}$ mapping inputs $\bm{\mathrm{x}}$ to a conditional distribution $p(\bm{\mathrm{C}}\mid\bm{\mathrm{x}})$ over the concepts, however it can be better understood as a function $\bm{\mathrm{\alpha}}:\mathcal{G}\to\Delta_{\mathcal{C}}$ taking ground-truth concepts $\bm{\mathrm{g}}$ as input instead, and defined as:
$\bm{\mathrm{\alpha}}(\bm{\mathrm{g}}):=\mathbb{E}_{\bm{\mathrm{x}}\sim p^{*}(\bm{\mathrm{X}}\mid\bm{\mathrm{g}})}[\bm{\mathrm{f}}(\bm{\mathrm{x}})]$ (2)
In contrast, the inference layer is a function $\bm{\mathrm{\omega}}:\Delta_{\mathcal{C}}\to\Delta_{\mathcal{Y}}$ mapping the concept distribution output by the concept extractor into a label distribution $p(\bm{\mathrm{Y}}\mid\bm{\mathrm{f}}(\bm{\mathrm{x}}))$. For clarity, we also define $\bm{\mathrm{\beta}}:\mathcal{C}\to\Delta_{\mathcal{Y}}$, which is identical to the former except it works with concept values rather than distributions, that is:
$\bm{\mathrm{\beta}}(\bm{\mathrm{c}}):=\bm{\mathrm{\omega}}(\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}\})$ (3)
Hence, a CBM entails both a pair $(\bm{\mathrm{f}},\bm{\mathrm{\omega}})\in\mathcal{F}\times\Omega$ and a pair $(\bm{\mathrm{\alpha}},\bm{\mathrm{\beta}})\in\mathcal{A}\times\mathcal{B}$. [^2] We work in the non-parametric setting, hence $\mathcal{F}$ and $\Omega$ contain all learnable concept extractors $\bm{\mathrm{f}}$ and inference layers $\bm{\mathrm{\omega}}$, and similarly $\mathcal{A}$ and $\mathcal{B}$ contain all learnable maps $\bm{\mathrm{\alpha}}$ and $\bm{\mathrm{\beta}}$. Later on, we will make use of the fact that $\mathcal{A}$ and $\mathcal{B}$ are simplices [60, 58], i.e., each $\bm{\mathrm{\alpha}}\in\mathcal{A}$ (resp. $\bm{\mathrm{\beta}}\in\mathcal{B}$) can be written as a convex combination of vertices $\mathsf{Vert}(\mathcal{A})$ (resp. $\mathsf{Vert}(\mathcal{B})$).
As mentioned in Section˜2, supplying prior knowledge $\mathsf{K}$ to a NeSy-CBM is equivalent to fixing the inference layer to a corresponding function $\bm{\mathrm{\omega}}^{*}$ (and $\bm{\mathrm{\beta}}^{*}$). Note that whereas $\bm{\mathrm{\omega}}^{*}$ changes based on how reasoning is implemented – e.g., fuzzy vs. probabilistic logic – $\bm{\mathrm{\beta}}^{*}$ does not, as both kinds of reasoning layers behave identically when input any point-mass concept distribution $\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}\}$.
Standard assumptions. The maps $\bm{\mathrm{\alpha}}$ and $\bm{\mathrm{\beta}}$, induced respectively by the concept extractor and inference layer, are especially useful for analysis provided the following two standard assumptions about how data are distributed [56, 91, 80]:
Assumption 3.1 (Extrapolability).
The ground-truth distribution $p^{*}(\bm{\mathrm{G}}\mid\bm{\mathrm{X}})$ is induced by a function $\bm{\mathrm{f}}^{*}:\bm{\mathrm{x}}\mapsto\bm{\mathrm{g}}$, i.e., $p^{*}(\bm{\mathrm{G}}\mid\bm{\mathrm{X}})=\mathbbm{1}\!\{\bm{\mathrm{G}}=\bm{\mathrm{f}}^{*}(\bm{\mathrm{X}})\}$.
This means that the ground-truth concepts $\bm{\mathrm{g}}$ can always be recovered for all inputs $\bm{\mathrm{x}}$ by a sufficiently expressive concept extractor; the $\bm{\mathrm{\alpha}}$ it induces is the identity $\mathrm{id}(\bm{\mathrm{g}}):=\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{g}}\}\in\mathsf{Vert}(\mathcal{A})$.
Assumption 3.2 (Deterministic knowledge).
The ground-truth distribution $p^{*}(\bm{\mathrm{Y}}\mid\bm{\mathrm{G}};\mathsf{K})$ is induced by the knowledge via a map $\bm{\mathrm{\beta}}^{*}\in\mathsf{Vert}(\mathcal{B})$, such that $p^{*}(\bm{\mathrm{Y}}\mid\bm{\mathrm{G}};\mathsf{K})=\bm{\mathrm{\beta}}^{*}(\bm{\mathrm{g}})$.
This ensures that the labels $\bm{\mathrm{y}}$ can be predicted without any ambiguity from the ground-truth concepts $\bm{\mathrm{g}}$. Notice that, not all choices of $\mathcal{G}$ guarantee that these assumptions are met. Small concept spaces may not give a deterministic knowledge, i.e., labels can be confused for one another with only few concepts, or too-arbitrary choices of the constituents may not guarantee their extrapolation from the input, i.e., ground-truth concepts of the input are ambiguous. Nonetheless, both assumptions hold in several NeSy tasks [4] and underlie many works in RSs [56, 53, 84], but formulating a theory that relaxes them is not straightfoward and is technically challenging. In fact, Marconato et al. [56] showed that upon relaxing Assumption˜3.2 it might not be possible to deal with RSs.
The maximum log-likelihood objective is then written as:
$\max_{(\bm{\mathrm{f}},\bm{\mathrm{\omega}})\in\mathcal{F}\times\Omega}\mathbb{E}_{(\bm{\mathrm{x}},\bm{\mathrm{y}})\sim p^{*}(\bm{\mathrm{X}},\bm{\mathrm{Y}})}\log(\bm{\mathrm{\omega}}_{\bm{\mathrm{y}}}\circ\bm{\mathrm{f}})(\bm{\mathrm{x}})$ (4)
Here, $\bm{\mathrm{\omega}}_{\bm{\mathrm{y}}}$ is the conditional probability of the ground-truth labels $\bm{\mathrm{y}}$. Notice that under the above assumptions CBMs attaining maximum likelihood perfectly model the ground-truth data distribution $p^{*}(\bm{\mathrm{X}},\bm{\mathrm{Y}})$, see Lemma˜10.1.
3.1 Intended Semantics and Joint Reasoning Shortcuts
We posit that a CBM $(\bm{\mathrm{\alpha}},\bm{\mathrm{\beta}})\in\mathcal{A}\times\mathcal{B}$ has “high-quality” concepts and inference layer if it satisfies two desiderata: (i) disentanglement: each learned concept $C_{i}$ should correspond to a single ground-truth concept $G_{j}$ up to an invertible transformation; (ii) generalization: the combination of $\bm{\mathrm{\alpha}}$ and $\bm{\mathrm{\beta}}$ must always yield correct predictions.
In our setting, without concept supervision and prior knowledge, the learned concepts are anonymous and users have to figure out which is which in a post-hoc fashion, e.g., by aligning them to dense annotations [39, 21, 13, 79]. Doing so is also a prerequisite for understanding the learned inference layer [87, 13, 21, 22]. When disentanglement holds the mapping between $\bm{\mathrm{G}}$ and $\bm{\mathrm{C}}$ can be recovered exactly using Hungarian matching [40]; otherwise it is arbitrarily difficult to recover, hindering interpretability. This links CBMs with many works that treat model representations’ identifiability in Independent Component Analysis and Causal Representation Learning [35, 27, 46, 41, 82], where disentanglement plays a central role. Generalization is equally important, as it entails that CBMs generalize beyond the support of training data and yields sensible OOD behavior, i.e., output the same predictions that would be obtained by using the ground-truth pair. Therefore, the pairs $(\bm{\mathrm{\alpha}},\bm{\mathrm{\beta}})$ which satisfy both desiderata will be equivalent to the ground-truth pair $(\mathrm{id},\bm{\mathrm{\beta}}^{*})$ and equally valid solutions for the NeSy task. Since we want $\bm{\mathrm{\alpha}}$ to be disentangled, this implies, in turn, a specific form for the map $\bm{\mathrm{\beta}}$, as shown by the next definition, which formalizes these desiderata:
Definition 3.3 (Intended Semantics).
A CBM $(\bm{\mathrm{f}},\bm{\mathrm{\omega}})\in\mathcal{F}\times\Omega$ entailing a pair $(\bm{\mathrm{\alpha}},\bm{\mathrm{\beta}})\in\mathcal{A}\times\mathcal{B}$ possesses the intended semantics if there exists a permutation $\pi:[k]\to[k]$ and $k$ element-wise invertible functions $\psi_{1},\ldots,\psi_{k}$ such that:
$\bm{\mathrm{\alpha}}(\bm{\mathrm{g}})=(\bm{\mathrm{\psi}}\circ\bm{\mathrm{P}}_{\pi}\circ\mathrm{id})(\bm{\mathrm{g}})\qquad\qquad\bm{\mathrm{\beta}}(\bm{\mathrm{c}})=(\bm{\mathrm{\beta}}^{*}\circ\bm{\mathrm{P}}_{\pi}^{-1}\circ\bm{\mathrm{\psi}}^{-1})(\bm{\mathrm{c}})$ (5)
Here, $\bm{\mathrm{P}}_{\pi}:\mathcal{C}\to\mathcal{C}$ is the permutation matrix induced by $\pi$ and $\bm{\mathrm{\psi}}(\bm{\mathrm{c}}):=(\psi_{1}(c_{1}),\ldots,\psi_{k}(c_{k}))$. In this case, we say that $(\bm{\mathrm{\alpha}},\bm{\mathrm{\beta}})$ is equivalent to $(\mathrm{id},\bm{\mathrm{\beta}}^{*})$, i.e., $(\bm{\mathrm{\alpha}},\bm{\mathrm{\beta}})\sim(\mathrm{id},\bm{\mathrm{\beta}}^{*})$
Intended semantics holds if the learned concepts $\bm{\mathrm{C}}$ match the ground-truth concepts $\bm{\mathrm{G}}$ modulo simple invertible transformations – like reordering and negation (Eq.˜5, left) – and the inference layer undoes these transformations before applying the “right” inference layer $\bm{\mathrm{\beta}}^{*}$ (Eq.˜5, right). In particular, Eq.˜5 (left) guarantees disentanglement of the concepts, ensuring that each learned concept corresponds to a distinct ground-truth concept up to an invertible transformation, matching the notion of [41]. A similar equivalence relation was analyzed for continuous representations in energy-based models, including supervised classifiers, by Khemakhem et al. [35]. CBMs satisfying these conditions are equivalent – specifically by the equivalence relation $\sim$, see Section˜10.1 – to the ground-truth pair $(\mathrm{id},\bm{\mathrm{\beta}}^{*})$; see Fig.˜2 (middle) for an illustration. In Lemma˜10.4, we prove that models with the intended semantics yield the same predictions of the ground-truth pair for all $\bm{\mathrm{g}}\in\mathcal{G}$.
Training a (NeSy) CBM via maximum likelihood does not guarantee it will embody intended semantics. We denote these failure cases as joint reasoning shortcuts (JRSs):
Definition 3.4 (Joint Reasoning Shortcut).
Take a CBM $(\bm{\mathrm{f}},\bm{\mathrm{\omega}})\in\mathcal{F}\times\Omega$ that attains optimal log-likelihood (Eq.˜4). The pair $(\bm{\mathrm{\alpha}},\bm{\mathrm{\beta}})$ entailed by it (Eqs.˜2 and 3) is a JRSs if it does not possess the intended semantics (Definition˜3.3), i.e., $(\bm{\mathrm{\alpha}},\bm{\mathrm{\beta}})\not\sim(\mathrm{id},\bm{\mathrm{\beta}}^{*})$.
JRSs can take different forms. First, even if the learned inference layer $\bm{\mathrm{\beta}}$ matches (modulo $\bm{\mathrm{P}}_{\pi}$ and $\bm{\mathrm{\psi}}$) the ground-truth one $\bm{\mathrm{\beta}}^{*}$, $\bm{\mathrm{\alpha}}$ might not match (also modulo $\bm{\mathrm{P}}_{\pi}$ and $\bm{\mathrm{\psi}}$) the identity. This is analogous to regular RSs in NeSy CBMs (Section˜2), in that the learned concepts do not reflect the ground-truth ones: while this is sufficient for high in-distribution performance (the training likelihood is in fact optimal), it may yield poor OOD behavior. Second, even if $\bm{\mathrm{\alpha}}$ matches the identity, the inference layer $\bm{\mathrm{\beta}}$ might not match the ground-truth one $\bm{\mathrm{\beta}}^{*}$. In our experiments Section˜5, we observe that this often stems from simplicity bias [92], i.e., the CBM’s inference layer tends to acquire specialized functions $\bm{\mathrm{\beta}}$ that are much simpler than $\bm{\mathrm{\beta}}^{*}$, leading to erroneous predictions OOD. Finally, neither the inference layer $\bm{\mathrm{\beta}}$ nor the map $\bm{\mathrm{\alpha}}$ might match the ground-truth, opening the door to additional failure modes. Consider the following example:
Example 3.5.
Consider a MNIST-SumParity problem where the training set consists of all possible unique (even, even), (odd, odd), and (odd, even) pairs of MNIST digits, as in Fig.˜1. A CBM trained on this data would achieve perfect accuracy by learning a JRS that extracts the parity of each input digit, yielding two binary concepts in $\{0,1\}$, and computes the difference between these two concepts. This JRS involves much simpler concepts and knowledge compared to the ground-truth ones. It also mispredicts all OOD pairs where the even digits come before odd digits.
While the presence of JRSs undermines both interpretability and OOD generalization, their absence alone does not guarantee OOD robustness. However, if JRSs are present, the likelihood of OOD failure increases substantially. For example, consider Example˜3.5, where a CBM achieves perfect training accuracy by learning a shortcut that captures only the parity of each digit rather than the full ground-truth concepts. This shortcut fails on OOD pairs; for example, given an (even, odd) pair not present in the training set, the model wrongly outputs 1-1 as the final label.
3.2 Theoretical Analysis
We now count the deterministic JRSs admitted by a task, i.e., those that lie on the vertices of $\mathcal{A}$ and $\mathcal{B}$: $(\bm{\mathrm{\alpha}},\bm{\mathrm{\beta}})\in\mathsf{Vert}(\mathcal{A})\times\mathsf{Vert}(\mathcal{B})$. We then show Theorem˜3.9 that this number determines whether general (non-deterministic) JRSs exist.
Theorem 3.6 (Informal).
Under Assumptions˜3.1 and 3.2, the number of deterministic JRSs is:
$\sum_{(\bm{\mathrm{\alpha}},\bm{\mathrm{\beta}})}\mathbbm{1}\!\{\bigwedge_{\bm{\mathrm{g}}\in\mathrm{supp}(\bm{\mathrm{G}})}(\bm{\mathrm{\beta}}\circ\bm{\mathrm{\alpha}})(\bm{\mathrm{g}})=\bm{\mathrm{\beta}}^{*}(\bm{\mathrm{g}})\}-C[\mathcal{G}]$ (6)
where the sum runs over $\mathsf{Vert}(\mathcal{A})\times\mathsf{Vert}(\mathcal{B})$, and $C[\mathcal{G}]$ counts the pairs with intended semantics.
All proofs and the definition of $C[\mathcal{G}]$ can be found in Section˜10. Intuitively, the first count includes all deterministic pairs $(\bm{\mathrm{\alpha}},\bm{\mathrm{\beta}})$ that achieve maximum likelihood on the training set (i.e., the predicted labels $(\bm{\mathrm{\beta}}\circ\bm{\mathrm{\alpha}})(\bm{\mathrm{g}})$ matches the ground-truth one $\bm{\mathrm{\beta}}^{*}(\bm{\mathrm{g}})$ for all ground-truth concepts $\bm{\mathrm{g}}$ appearing in the data), while the second term subtracts those $(\bm{\mathrm{\alpha}},\bm{\mathrm{\beta}})$ that possess the intended semantics as per Definition˜3.3. A positive count implies that a CBM trained via maximum likelihood can learn a deterministic JRS. Notice that the count of JRSs also depends on the choice of the concept space $\mathcal{G}$: a large number of fine-grained ground-truth concept may increase the number of JRSs, while coarser-grained concepts that are lower in number may reduce them.
As a sanity check, we show that if prior knowledge $\mathsf{K}$ is provided – fixing the inference layer to $\bm{\mathrm{\beta}}^{*}$ – the number of deterministic JRSs in fact matches that of RSs, as expected:
Corollary 3.7.
Consider a fixed $\bm{\mathrm{\beta}}^{*}\in\mathsf{Vert}(\mathcal{B})$. Under Assumptions˜3.1 and 3.2, the number of deterministic JRSs $(\bm{\mathrm{\alpha}},\bm{\mathrm{\beta}}^{*})\in\mathsf{Vert}(\mathcal{A})\times\mathsf{Vert}(\mathcal{B})$ is:
$\sum_{\bm{\mathrm{\alpha}}\in\mathsf{Vert}(\mathcal{A})}\mathbbm{1}\!\{\bigwedge_{\bm{\mathrm{g}}\in\mathrm{supp}(\bm{\mathrm{G}})}(\bm{\mathrm{\beta}}^{*}\circ\bm{\mathrm{\alpha}})(\bm{\mathrm{g}})=\bm{\mathrm{\beta}}^{*}(\bm{\mathrm{g}})\}\textstyle-1$ (7)
This matches the count for deterministic RSs in [56].
Corollary˜3.7 implies that when $|\mathsf{Vert}(\mathcal{B})|>1$, the number of deterministic JRSs in Eq.˜6 can be much larger than the number of deterministic RSs (Eq.˜7). For example, in MNIST-SumParity with digits restricted to the range $[0,4]$ there exist $64$ RSs but about $100$ thousand JRSs, and the gap increases as we extend the range $[0,N]$. An in-depth analysis appears in Section˜9.5.
Next, we show that whenever the number of deterministic JRSs in Eq.˜6 is zero, there exist no JRSs at all, including non-deterministic ones. This however only applies to CBMs that satisfy the following natural assumption:
Assumption 3.8 (Extremality).
The inference layer $\bm{\mathrm{\omega}}\in\Omega$ satisfies extremality if, for all $\lambda\in(0,1)$ and for all $\bm{\mathrm{c}}\neq\bm{\mathrm{c}}^{\prime}\in\mathcal{C}$ such that $\operatorname*{argmax}_{\bm{\mathrm{y}}\in\mathcal{Y}}\bm{\mathrm{\omega}}(\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}\})_{\bm{\mathrm{y}}}\neq\operatorname*{argmax}_{\bm{\mathrm{y}}\in\mathcal{Y}}\bm{\mathrm{\omega}}(\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}^{\prime}\})_{\bm{\mathrm{y}}}$, it holds:
$\max_{\bm{\mathrm{y}}\in\mathcal{Y}}\bm{\mathrm{\omega}}(\lambda\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}\}+(1-\lambda)\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}^{\prime}\})_{\bm{\mathrm{y}}}<\max\left(\max_{\bm{\mathrm{y}}\in\mathcal{Y}}\bm{\mathrm{\omega}}(\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}\})_{\bm{\mathrm{y}}},\max_{\bm{\mathrm{y}}\in\mathcal{Y}}\bm{\mathrm{\omega}}(\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}^{\prime}\})_{\bm{\mathrm{y}}}\right)$
Intuitively, a CBM satisfies this assumption if its inference layer $\bm{\mathrm{\omega}}$ is “peaked”: for any two concept vectors $\bm{\mathrm{c}}$ and $\bm{\mathrm{c}}^{\prime}$ yielding distinct predictions, the label probability output by $\bm{\mathrm{\omega}}$ for any mixture distribution thereof is no larger than the label probability that it associates to $\bm{\mathrm{c}}$ and $\bm{\mathrm{c}}^{\prime}$. That is, distributing probability mass across concepts does not increase label likelihood. While this assumption does not hold for general CBMs, we show in Section˜12 that it holds for popular architectures, including most of those that we experiment with. Under Assumption˜3.8, we have:
Theorem 3.9 (Identifiability).
Under Assumptions˜3.1 and 3.2, if the number of deterministic JRSs (Eq.˜6) is zero then every CBM $(\bm{\mathrm{f}},\bm{\mathrm{\omega}})\in\mathcal{F}\times\Omega$ satisfying Assumption˜3.8 that attains maximum log-likelihood (Eq.˜4) possesses the intended semantics (Definition˜3.3). That is, the pair $(\bm{\mathrm{\alpha}},\bm{\mathrm{\beta}})\in\mathcal{A}\times\mathcal{B}$ entailed by each such CBM is equivalent to the ground-truth pair $(\mathrm{id},\bm{\mathrm{\beta}}^{*})$, i.e., $(\bm{\mathrm{\alpha}},\bm{\mathrm{\beta}})\sim(\mathrm{id},\bm{\mathrm{\beta}}^{*})$.
Clearly, a similar conclusion does not hold for models that do not satisfy Assumption˜3.8: even when deterministic JRSs are absent, a CBM can still be affected by non-deterministic JRSs.
4 Practical Solutions
By Theorem˜3.9, getting rid of deterministic JRSs from a task prevents CBMs trained via maximum log-likelihood to learn JRS altogether. The key question is how to make the JRSs count be zero. As previously explored in RSs, pairing the maximum likelihood objective with other well-known penalties in the literature can limit optimal $\bm{\mathrm{\alpha}}$’s for the joint training objective. Several mitigation strategies for RSs have shown to decrease the count of the $\bm{\mathrm{\alpha}}$’s, in some cases leading to zeroing deterministic RSs [56, 91]. We report the count reduction of different mitigations in Section˜11.
Supervised strategies. The most direct strategies for controlling the semantics of learned CBMs rely on supervision. Concept supervision is the go-to solution for improving concept quality in CBMs [39, 12, 21, 54] and NeSy-CBMs [4, 91]. Full concept supervision prevents learning $\bm{\mathrm{\alpha}}\neq\mathrm{id}$. However, this does not translate into guarantees on the inference layer $\bm{\mathrm{\beta}}$, at least in tasks affected by confounding factors such as spurious correlations among concepts [stammer2021right]. E.g., if the $i$-th and $j$-th ground-truth concepts are strongly correlated, the inference layer $\bm{\mathrm{\beta}}$ can interchangeably use either, regardless of what $\bm{\mathrm{\beta}}^{*}$ does.
Another option is employing knowledge distillation [6], that is, supplying supervision of the form $(\bm{\mathrm{g}},\bm{\mathrm{y}})$ to encourage the learned knowledge $\bm{\mathrm{\beta}}$ to be close to $\bm{\mathrm{\beta}}^{*}$ for the supervised $\bm{\mathrm{g}}$’s. This supervision is available for free provided concept supervision is available, but can also be obtained separately, e.g., by interactively collecting user interventions [72, 22, 76]. This strategy cannot avoid all JRSs because even if $\bm{\mathrm{\beta}}=\bm{\mathrm{\beta}}^{*}$, the CBM may suffer from regular RSs (i.e., $\bm{\mathrm{\alpha}}$ does not match the identity function $\mathrm{id}$).
Another option is to fit a CBM on multiple tasks [7] sharing concepts. It is known that a NeSy-CBM attaining optimal likelihood on multiple NeSy tasks with different prior knowledges is also optimal for the single NeSy task obtained by conjoining those knowledges [56]: this more constrained knowledge better steers the semantics of the learned concepts, ruling out potential JRSs. Naturally, collecting a sufficiently large number of diverse tasks using the same concepts can be highly non-trivial, depending on the application.
Unsupervised strategies. Many popular strategies for improving concept quality in (NeSy) CBMs are unsupervised. A cheap but effective one when dealing with multiple inputs is to process inputs individually, preventing mixing between their concepts. E.g., In MNIST-SumParity one can apply the same digit extractor to each digit separately, while for images with multiple objects one can first segment them (e.g., with YoloV11 [32]) and then process the resulting bounding boxes individually. This is extremely helpful for reducing, and possibly overcoming, RSs [56] and frequently used in practice [50, 81, 13, 79]. We apply it in all our experiments.
Both supervised [54] and unsupervised [alvarez2018towards, 44] CBMs may employ a reconstruction penalty [2, 37] to encourage learning informative concepts. Reconstruction can help prevent collapsing distinct ground-truth concepts into single learned concepts, e.g., in MNIST-SumParity the odd digits cannot be collapsed together without impairing reconstruction. [^3] This is provably the case context-style separation i.e., assuming concepts are independent of stylistic features like calligraphy [56, Proposition 6]. Alternatively, one can employ entropy maximization [51] to spread concept activations evenly across the bottleneck. This can be viewed as a less targeted but more efficient alternative to reconstruction, which becomes impractical for higher dimensional inputs. Another option is contrastive learning [11], in that augmentations render learned concepts more robust to style variations [83], e.g., for MNIST-based tasks it helps to cluster distinct digits [68]. Unsupervised strategies all implicitly counteract simplicity bias whereby $\bm{\mathrm{\alpha}}$ ends up collapsing.
5 Case Studies
Table 1: Results for MNIST-Add.
Method $F_{1}(Y)$ ($\uparrow$) $F_{1}(C)$ ($\uparrow$) $\mathsf{Cls}(C)$ ($\downarrow$) $F_{1}(\bm{\mathrm{\beta}})$ ($\uparrow$)
DPL $8\pm 0.01 0.99±0$$9\pm 0.01$ $0.01\pm 0.01$ -
[…]
SENN $0.97\pm 0.01$ $80\pm 0.07 0.0$$01\pm 0.01$ $0.75\pm 0.08$
[…]
DPL $0.91\pm 0.09$ $92\pm 0.08 0.0$$01\pm 0.01$ $0.90\pm 0.08$
[…]
Table 2: Results for MNIST-SumParity.
Method $F_{1}(Y)$ ($\uparrow$) $F_{1}(C)$ ($\uparrow$) $\mathsf{Cls}(C)$ ($\downarrow$) $F_{1}(\bm{\mathrm{\beta}})$ ($\uparrow$)
DPL $0.99\pm 0.01$ $0.43\pm 0.08$ $0.36\pm 0.15$ -
[…]
SENN $0.99\pm 0.01$ $49\pm 0.05 0.0$$01\pm 0.01$ $0.53\pm 0.06$
[…]
DPL $99\pm 0.01 0.0$$07\pm 0.01$ $0.80\pm 0.01$ $0.50\pm 0.05$
[…]
Table 3: Results for Clevr.
Method $F_{1}(Y)$ ($\uparrow$) $F_{1}(C)$ ($\uparrow$) $\mathsf{Cls}(C)$ ($\downarrow$) $F_{1}(\bm{\mathrm{\beta}})$ ($\uparrow$)
DPL $0.99\pm 0.01$ $0.25\pm 0.05$ $0.57\pm 0.02$ -
[…]
DPL $0.99\pm 0.01$ $0.22\pm 0.04$ $57\pm 0.03 0.0$$01\pm 0.01$
Table 4: CBNM on biased MNIST-SumParity.
ID OOD
C-Sup K-Sup $F_{1}(Y)$ $(\uparrow)$ $F_{1}(\bm{\mathrm{\beta}})$ $(\uparrow)$ $F_{1}(Y)$ $(\uparrow)$ $F_{1}(\bm{\mathrm{\beta}})$ $(\uparrow)$
$\% 00\%$ $0.99\pm 0.01$ $55\pm 0.05 0.0$$01\pm 0.01$ $0.47\pm 0.07$
[…]
$100\%$ $0\%$ $0.97\pm 0.01$ $0.88\pm 0.04$ $0.40\pm 0.14$ $0.31\pm 0.12$
[…]
Table 5: Results for BDD-OIA.
Method $F_{1}(Y)$ ($\uparrow$) $F_{1}(C)$ ($\uparrow$) $\mathsf{Cls}(C)$ ($\downarrow$) $F_{1}(\bm{\mathrm{\beta}})$ ($\uparrow$)
DPL $0.69\pm 0.03$ $0.44\pm 0.01$ $0.88\pm 0.01$ -
[…]
We tackle the following key research questions: Q1. Are CBMs affected by JRSs in practice? Q2. Do JRSs affect interpretability and OOD behavior? Q3. Can existing mitigation strategies prevent JRSs? Section˜8 reports additional details about the tasks, architectures, and model selection.
Models. We evaluate several (also NeSy) CBMs. DeepProbLog (DPL) [50, 51] is the only method supplied with the ground-truth knowledge $\mathsf{K}$, and uses probabilistic-logic reasoning to ensure predictions are consistent with it. CBNM is a Concept-Bottleneck Model [39] with no supervision on the concepts. Deep Symbolic Learning (DSL) [13] is a SOTA NeSy-CBM that learns concepts and symbolic knowledge jointly; it implements the latter as a truth table and reasoning using fuzzy logic. We also consider two variants: DPL replaces fuzzy with probabilistic logic, while bears wraps DPL within BEARS [53], an ensemble approach for handling regular RSs. In short, bears consists of a learned inference layer and multiple concept extractors. The former is learned only once, the latter are learned sequentially. We also evaluate Self-explainable Neural Networks (SENN) [alvarez2018towards], unsupervised CBMs that include a reconstruction penalty. Since they do not satisfy our assumptions, they are useful to empirically assess the generality of our remarks. In our experiments, we train all CBM variants in a joint manner [39] using only label supervision and do not rely on foundation models for gathering annotations [61, 93, 74]. For additional details, see Section˜8.
Data sets. To assess both learned concepts and inference layer, we use three representative NeSy tasks with explicit concept annotations and prior knowledge. MNIST-Add [50] involves predicting the sum of two MNIST digits, and serves as a sanity check as it provably contains no RSs. MNIST-SumParity is similar except we have to predict the parity of the sum, as in Fig.˜1, and admits many JRSs. We also include a variant of Clevr [33] in which images belong to $3$ classes as determined by a logic formula and only contain $2$ to $4$ objects, for scalability. Finally, we carry out basic checks also on BDD-OIA [90], a real-world autonomous driving task where images are annotated with actions (move_forward, stop, turn_left, turn_right) and 21 binary concepts.
Metrics. For each CBM, we evaluate predicted labels and concepts with the $F_{1}$ score (resp. $F_{1}(Y)$ and $F_{1}(C)$) on the test split. Computing $F_{1}(C)$ requires to first align the predicted concepts to ground-truth annotations. We do so by using the Hungarian algorithm [40] to find a permutation $\pi$ that maximizes the Pearson correlation across concepts. In this way, we directly evaluate disentanglement of the concepts as per Eq.˜5 (left). Concept collapse $\mathsf{Cls}(C)$ quantifies to what extent the concept extractor blends different ground-truth concepts together: high collapse indicates that it predicts fewer concepts than expected. We measure the quality of the learned inference layer $\bm{\mathrm{\beta}}$ as follows: for each input $\bm{\mathrm{x}}$, (1) we reorder the corresponding concept annotations $\bm{\mathrm{g}}$ using $\pi$, (2) feed the result to $\bm{\mathrm{\beta}}$, and (3) compute the $F_{1}$-score ($F_{1}(\bm{\mathrm{\beta}})$) of its predictions. Step (1) permutes and applies element-wise invertible transformations on the ground-truth concepts using the inverse of the map returned by Hungarian matching. This evaluates whether the learned inference layer can correctly handle (an invertible transformation of) the ground-truth concept annotations, measuring the generalization as per Eq.˜5 (right).This procedure is precisely detailed in Section˜8.6.
Mitigation strategies. We evaluate all strategies discussed in Section˜4 as well as a combination of all unsupervised mitigations, denoted H+R+C (entropy, reconstruction, contrastive). Implementations are taken verbatim from [4]. For contrastive learning, we apply an InfoNCE loss to the predicted concept distribution with the augmentations suggested by Chen et al. [11]. For knowledge distillation, the inference layer is fed ground-truth concepts and trained to optimize the log-likelihood of the corresponding labels.
Q1: All tested CBMs suffer from JRSs and simplicity bias. We analyze three learning tasks of increasing complexity. MNIST-Add provably contains no JRSs and thus satisfies the preconditions of Theorem˜3.9. Compatibly, all CBMs attain high label performance ($F_{1}(Y)\geq 90\%$) and high-quality semantics, as shown by the values of $F_{1}(C)$ ($\geq 90\%$) and $F_{1}(\bm{\mathrm{\beta}})$ ($\geq 75\%$) seen in Table˜5. The only exception is bears, which by design trades off performance for calibration. However, in MNIST-SumParity and Clevr, [^4] For Clevr, we focus on representative CBMs with high fit on the training data. which are more complex, all models lack intended semantics. This is clear from Tables˜5 and 5: despite attaining high $F_{1}(Y)$ ($\geq 90\%$), the learned concepts and inference layer are low quality: $F_{1}(C)$ is always low – it never crosses the $50\%$ mark – and $F_{1}(\bm{\mathrm{\beta}})$ is at best around chance level. Our results on BDD-OIA in Table˜5 show the very same trend.
The behavior of concept collapse strongly hints at simplicity bias. While no collapse takes place in MNIST-Add ($\mathsf{Cls}(C)\leq 0.01$, i.e., digits are not compacted together), in MNIST-SumParity and Clevr collapse is omnipresent ($\mathsf{Cls}(C)\geq 0.22$), suggesting that CBMs are prone to simplicity bias, as expected. SENN is an outlier, likely due to the higher capacity of its inference layer (rather than reconstruction, which is entirely ineffective in Q3). Even having access to the ground-truth knowledge is ineffective, as shown by the high degree of collapse displayed by DPL.
Q2: JRSs compromise OOD behavior and supervision only partially helps. We evaluate the impact of JRSs and supervision on OOD behavior by training a CBNM on the biased version of MNIST-SumParity in Fig.˜1. The in-distribution (ID) data comprise all (odd, odd), (even, even), and (odd, even) pairs of MNIST digits, while the OOD split contains all (even, odd) pairs. Table˜5 reports performance under varying levels of concept and knowledge supervision. While all models produce high-quality predictions in-distribution ($F_{1}(Y)\geq 0.97$), only the CBNM receiving complete supervision attains acceptable OOD predictions: for the others, $F_{1}(Y)\leq 0.40$. Even in this case, though, the inference layer still does not have intended semantics, as shown by $F_{1}(\bm{\mathrm{\beta}})\leq 70\%$.
Q3: Popular CBM and RS mitigations fall short. Finally, we test mitigation strategies on CBNM trained on MNIST-SumParity and Clevr, ablating the amount of supervision. The results for MNIST-SumParity in Fig.˜3 show while concept supervision ($x$-axis) helps all metrics, adding reconstruction, entropy maximization, and contrastive learning to the mix brings no benefits for concept and knowledge ($F_{1}(C)$, $F_{1}(\bm{\mathrm{\beta}})$) quality, not even when combined (orange curve). Knowledge supervision is also ineffective (blue lines), likely because MNIST-SumParity and Clevr suffer from RSs even when the model is supplied prior knowledge. Contrastive learning improves concept collapse ($\mathsf{Cls}(C)$): the orange and purple curves in Fig.˜3 (right) show it can prevent CBMs from mixing concepts together. The results for Clevr in Section˜9 show similar trends.
Figure 3: Traditional mitigations have limited effect on CBNM for MNIST-SumParity. The only outlier is contrastive learning (orange and purple), which consistently ameliorates concept collapse.
6 Related Work
Shortcuts in machine learning. Shortcuts occur when machine learning models solve a prediction task by using features that correlate with but do not cause the desired output, leading to poor OOD behavior [geirhos2020shortcut, teso2023leveraging, 95, 75]. Existing work focuses on black-box models rather than CBMs. Existing studies on the semantics of CBM concepts [57, 49, 29, 65] focus on individual failures but lack formal notion of intended semantics. One exception is [55] which, however, ignores the role of the inference layer. We build on known results on reasoning shortcuts [45, 56, 84, 80] which are restricted to NeSy-CBMs in which the prior knowledge is given and fixed. Our analysis upgrades these results to general CBMs. Furthermore, our simulations indicate that strategies that are effective for CBMs and RSs no longer work for JRSs. At the same time, while NeSy approaches that learn prior knowledge and concepts jointly are increasingly popular [85, 94, 48, 13, 79, 87], existing work neglects the issue of shortcuts in this setting. Our work partially fills this gap. Shortcuts and JRSs are related but distinct. While shortcuts can induce JRSs, the converse is not true: a model can be affected by a JRS even if it does not rely on confounders, as shown in Fig.˜2. This distinction also implies that mitigation strategies designed for vanilla shortcuts do not necessarily transfer to JRSs.
Relation to identifiability. Works in Independent Component Analysis and Causal Representation Learning focus on the identifiability of representations up to an equivalence relation, typically up to permutations and rescaling [34, 26, 82] or more general transformations [67, 5]. The equivalence relation introduced along with intended semantics (Definition˜3.3) establishes a specific connection between the maps $\bm{\mathrm{\alpha}}$ and $\bm{\mathrm{\beta}}$. This aligns with existing works that aim to identify representations and linear inference layers using supervised labels [41, 23, 3, nielsen2024challenges, nielsen2025does]. Moreover, while prior works primarily focus on identifying real-valued representations, we examine the identifiability of discrete concepts and of the inference layer. [64] explores the identifiability of (potentially discrete) concepts linearly encoded in model representations but focuses on multi-environment settings. Our results differ in that we show how concepts can be identified in CBMs by circumventing reasoning shortcuts. A concurrent work by Goyal et al. [25] identifies latent concepts through label supervision and a sparsity regularization, showing that a variant of CBNMs can provably recover concepts with the intended semantics and thus avoiding JRSs altogether, see [25, Theorem 2].
7 Conclusion
We study the issue of learning CBMs and NeSy-CBMs with high-quality concepts and inference layers. We formalize intended semantics and joint reasoning shortcuts (JRSs), showing how, under suitable assumptions, zeroing the number of deterministic JRSs provably prevents all JRSs, yielding identifiability. Numerical simulations indicate that JRSs are very frequent and impactful, and that the only (partially) beneficial mitigation stratgy is contrastive learning. Our work paves the way to the design of effective solutions and therefore more robust and interpretable CBMs. In future work, we plan to extend our theory to general neural nets and large language models, and we will take a closer look at more complex mitigations such as smartly constraining the inference layer [48, 87], debiasing [bahadori2021debiasing], and human interaction [teso2023leveraging]. Another asset is considering concept-level supervision gathered from foundation models like CLIP, however recent results [15] call for caution in using them to supervise the bottleneck. Wrong concept-level supervision may transfer biases of foundation models and induce JRSs, thus resulting ineffective as a mitigation strategy.
Broader impact. With this work, we aim to highlight the issue of joint reasoning shortcuts in concept-based models and their subtle but impactful consequences on the trustworthiness and interpretability of these models, as well as on their reliability in out-of-distribution scenarios. It also highlights the limited effect of unsupervised mitigation strategies, thus pointing out a significant gap in our toolbox for addressing joint reasoning shortcuts effectively and cheaply. Our work also provides a solid theoretical foundation upon which further studies can build.
Limitations. Our theoretical analysis of JRSs builds on two common assumptions for the data generation process, which limit its scope. Relaxing them, while not straightforward, will allow treating models and NeSy tasks with intrinsic uncertainty on label and concepts. Our experiments span both synthetically-generated and a real-world vision datasets; further analysis should highlight the impact of JRSs in other real-world settings and beyond the vision domain.
Acknowledgements
The authors are grateful to Antonio Vergari, Emile van Krieken, Marco Pacini, and Luigi Gresele for useful discussions. Funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Health and Digital Executive Agency (HaDEA). Neither the European Union nor the granting authority can be held responsible for them. Grant Agreement no. 101120763 - TANGO. The work of AP was partially supported by the MUR PNRR project FAIR - Future AI Research (PE00000013) funded by the NextGenerationEU. PM was supported by the MSCA project “Probabilistic Formal Verification for Provably Trustworthy AI - PFV-4-PTAI” under GA no. 101110960.